# Domain Transfer Through Image-to-Image Translation for Uncertainty-Aware Prostate Cancer Classification

### v1

@@ 2.2 Deep Learning for PCa Classification
[…]
In Yoo et al. (2019), a probabilistic approach using mp-MRI data is employed for PCa classification. The authors develop an automated pipeline for the classification of clinically significant PCa using 3.0T DWI images from 427 patients. The pipeline consists of three parts: classification of each DWI slice using the pre-activated ResNet model (He et al., 2016), extraction and selection of first-order statistics from the CNN outputs, and final class label prediction using a random forest classifier. On the test set, the model achieves an AUC of 0.87. While the aforementioned studies may yield favorable AUCs, the reproducibility of the model might be challenging clinics with limited patient (data) throughput. Recently, Grebenisan et al. (2021, 2020) address the data-hungry problem by introducing a disentangled representation learning approach (SDNet) to synthesize public 3.0T MRI images into 1.5T MRI images to increase the training data size for centres with limited 1.5T data. Their approach aims to separate the anatomy- and modality-specific features present in images, subsequently merging the 1.5T modality features with the 3.0T anatomical features to generate MRI images resembling those acquired at 1.5T. Finally, a simple 3D-CNN classifier is used for the binary classification of clinically significant PCa. The model outperforms the state-of-the-art performance in PCa classification through domain alignment between different data sources.
Although current methods for PCa classification can achieve good performance, they do not provide a confidence score for their prediction, making them less interpretable in clinical practice.
@@ 2.3 Uncertainty Estimation
Recent studies in medical imaging have highlighted the detrimental impact of label noise on the performance of modern deep learning models (Karimi et al., 2020). Conventional regularization techniques such as dropout, batch normalization, weight decay, etc. can not properly address the problem (Arpit et al., 2017; Zhang et al., 2021). Methods proposed to mitigate such problem can be summarized as those that use (Song et al., 2022): (1) Robust loss functions and loss adjustments (Van Rooyen et al., 2015; Charoenphakdee et al., 2019; Zhang and Sabuncu, 2018) aiming to stabilize the model performance when optimizing its parameters; (2) Sample selection (Jiang et al., 2018; Malach and Shalev-Shwartz, 2017; Wang et al., 2020) aiming to select a subset of “clean” data from a batch of samples to compute the loss; and (3) Robust architectures (Han et al., 2018; Yu et al., 2019) aiming to learn the same data by training multiple models with different initialization assess output stability. While these methods inherently handle the noisy label problem, they can not provide explicit uncertainty estimation in terms of confidence in their output. Moreover, the capability of deep learning models to effectively identify irrelevant samples is still limited. For instance, when a model trained on prostate MRIs is presented with a CT scan of the prostate at the time of inference, it is unclear whether the model can provide meaningful predictions or simply indicate a lack of in-domain knowledge and perform a human-in-the-loop analysis instead. In recent years, research has been conducted on uncertainty estimation for deep learning models. Gal and Ghahramani (2016, 2015) develop the dropout neural networks framework to represent the prediction uncertainty of deep learning models, where the dropout layers in the model are formed by Bernoulli distributed random variables. During the test phase, the predictive uncertainty can be determined by enabling dropout layers and averaging the results of multiple runs. An alternative approach to modeling uncertainty involves the use of evidential neural networks (Sensoy et al., 2018), which formulate uncertainty by fitting a Dirichlet distribution - acting as the conjugate prior of the categorical distribution — to the class probabilities acquired from neural networks (Sensoy et al., 2018). This method considers model predictions as multinomial subjective opinions (Jøsang, 2016) or beliefs (Dempster, 1968), which can be further modeled explicitly using subjective logic. The "evidential" approach emphasizes the ability of the model to deliver certain predictions and exhibits superiority compared to the dropout approach (Gal and Ghahramani, 2016).
@@
In clinical practice, uncertainty estimation is crucial. By integrating uncertainty information into prediction outcomes, misclassification rates can be significantly reduced. For instance, in radiograph classification task (Ghesu et al., 2019), the authors employ the Dempster-Shafer Theory of Evidence (Dempster, 1968) and the principles of subjective logic (Jøsang, 2016) to develop a framework that jointly estimates per-class probabilities and provides predictive uncertainty. Later, this approach has been extended to abdominal ultrasound and brain MR images (Ghesu et al., 2021). In the context of breast cancer classification, Tardy et al. (2019) apply the evidential neural networks approach (Sensoy et al., 2018) to effectively diagnose breast cancer. A similar approach is used for the same task by Yuan et al. (2020) through the evidence adjustment technique, which focuses on the difference in the risks of uncertain samples from different classes. Consequently, we build upon the work from Sensoy et al. (2018) by adding uncertainty estimation to improve the robustness of the model and the interpretability of predictions.
@@ 3.1 Data
[…]
@@
Figure 1: Visualization of sample data. 0(a) and 0(b) are the 1.5T T2 and ADC images from KHSC, respectively. Similarly, 0(c) and 0(d) are the 3.0T T2 and ADC images from the “ProstateX” Challenge, respectively.
@@ 3.2 Pre-processing
T2 and ADC sequences from both datasets are $160\times 160\times C$, where $C$ is the total number of slices in the MRI. We resample all 3D data to have the same voxel spacing. To reduce aliasing artifacts, the most common voxel spacing ($0.5\times 0.5\times 3\leavevmode\nobreak\ mm^{3}$) is used across all data, and the consine-windowed interpolation is utilized during sampling. We normalize pixel intensities to $[-1,1]$ for all data. For the translation purpose from 3.0T to 1.5T, we further resample all 3D data to $256\times 256\times C$ and split into $C$ 2D gray-scale slices.
@@ 4.1 The Domain Transfer Framework
We adapt the ACL-GAN model (Zhao et al., 2020) discussed in Section 2.1 to perform unpaired MR image translation from 3.0T to 1.5T. There are two generators in this model namely $G_{T->S}$ and $G_{S->T}$, $G_{T->S}$ translates the images from the target domain to the source domain given the input $x$ and a noise vector $z$ sampled from $\mathcal{N}(0,1)$. $G_{S->T}$ is the reverse process of $G_{T->S}$ which translates the image from the source domain to the target domain. There are three discriminators, $D_{S},D_{T}$, and $\hat{D}$ in this model. The first two ensure that translated images are in the correct domain by optimizing adversarial losses, and $\hat{D}$ ensures that translated images retain anatomical features in 3.0T by distinguishing the pair (Source, Trans. Source1) and (Source, Trans. Source2), as shown in the bottom of Figure 2. The loss function of ACL-GAN (Zhao et al., 2020) is defined as in equation (1):
@@
Where $\mathcal{L}_{adv}$ is the traditional adversarial loss for both source domain $S$ and target domain $T$, i.e., $\mathcal{L}_{adv}=\mathcal{L}^{S}_{adv}+\mathcal{L}^{T}_{adv}$, to ensure the translated image is in the correct domain. $\mathcal{L}_{acl}$ is the adversarial consistency loss that is used to preserve important features of the source image in the translated image, $\mathcal{L}_{idt}$ is the identity loss, which encourages the generators to perform approximately identity mapping when images in the corresponding domain are provided, and $\mathcal{L}_{mask}$ is used to force both generators to only modify certain regions of the source image and keep the rest of the areas unchanged. Readers are encouraged to refer to the original paper (Zhao et al., 2020) for more details.
@@ 4.2.1 Classifier architectures
The traditional CNN approach is used for the clinically significant PCa binary classification task. Specifically, we explore three different model architectures for combinations of T2 and ADC patches. The first architecture, called the multi-stream CNN (“M.S. MpMRI”), treats T2 and ADC patches as separate inputs, as shown in Figure 3. The model takes 3D patches of T2 and ADC as parallel inputs, which are then processed by the same feature extractor to extract deep semantic representations. The output representations of T2 and ADC are then concatenated channel-wise and fed into another convolutional layer followed by a fully-connected layer to produce the class probabilities.
In the second architecture, we combine ADC and T2 patches as a single input to the network. We stack cropped 3D patches of T2 and ADC along the channel axis and obtain the input data size of $64\times 64\times 6$. Another way to combine is we only consider the located suspicious slice $i_{s}$ for both T2 and ADC, and stack them along the channel axis to obtain the input data size of $64\times 64\times 2$. The model architecture for both combinations is similar to Figure 3, where there is only one branch and no concatenation afterward. We name the model with input size of $64\times 64\times 6$ (resp. input size $64\times 64\times 2$) as “Vol. MpMRI”(resp. “MpMRI”).
Lastly, we use only 3D T2 patches as input to match with the previous work (Grebenisan et al., 2021). The model architecture is as same as the one for MpMRI, and we call this model “T2-only”.
@@
To combat the potential noisy label problem, the co-teaching framework (Han et al., 2018) is also utilized in this work. In co-teaching, two models with the same architecture and configurations are trained simultaneously, as shown in Figure 4. In every mini-batch, two models are trained in parallel. Each model first feeds forward all data in the current batch and selects some of the data that has the clean labels with high probability; then, two models decide on what data in the current batch should be used for training; finally, each model uses the data selected by the peer model to update itself. Denote $f$ be the first model and $g$ be the second model. The number of instances selected by both models is controlled by a non-increasing function $R(T)$ on $T$ defined in equation (2). At epoch $T$, each model only calculate loss based on the $R(T)$ portion of the batch instances.
Figure 4: The co-teaching framework, where A and B are two models with the same architectures that are trained in parallel and simultaneously.
$R_{T}=(1-\tau\cdot\min(T/T_{k},1))\times 100\%$ (2)
where $T$ is the total training epochs, $\tau$ is the same as noise rate, and $T_{k}$ is the epochs for linear drop rate. We use the “MpMRI” as the backbone model (model A and B in Figure 4) in the co-teaching framework.
@@ 4.2.2 Evidential Focal Loss
Dataset filtering during the training phase could reduce the effect of the noisy label on the deep model. Followed by Ghesu et al. (2019), the process of uncertainty-based filtering is shown at the top of Figure 2: Firstly, we calculate the uncertainty value for each sample in the training set. We then remove a portion of the training samples that exhibit high predictive uncertainty. Finally, we retrain the model using the remaining “clean” training data.
Following the work from Sensoy et al. (2018), we extend and combine the idea of subjective logic (Jøsang, 2016) with the focal loss (Lin et al., 2017) for the clinically significant PCa binary classification task. In the context of the Theory of Evidence, a belief mass is assigned to individual attributes, e.g., the possible class label of a specific data sample. The belief mass is generally calculated from the evidence collected from the observed data (Dempster, 1968). Let $K$ be the number of classes, and $b_{k}\geq 0,k\in[1,K]$ is the belief mass for class $k$ and $u\geq 0$ is the overall uncertainty measure. Let $e_{k}\geq 0$ be the evidence computed for $k^{th}$ class, then followed by Sensoy et al. (2018), the belief $b_{k}$ and the uncertainty $u$ are computed as the following:
@@
where $S=\sum_{i=1}^{K}(e_{i}+1)$. For our binary task ($K=2$), we can further simplify Equation (3) to $b_{0}=\frac{e_{0}}{e_{0}+e_{1}+2},b_{1}=\frac{e_{1}}{e_{0}+e_{1}+2}$, and $u=\frac{2}{e_{0}+e_{1}+2}$. The belief mass assignment, e.g., subjective opinion, is corresponding to the Dirichlet distribution with parameters $\alpha_{k}=e_{k}+1$, and we could rewrite $S=\sum_{i=1}^{K}\alpha_{k}$ as the Dirichlet strength. The formal definition of Dirichlet distribution can be found in Sensoy et al. (2018). The expected probability for the $k^{th}$ class is calculated by the mean with the associate Dirichlet distribution $\hat{p}_{k}=\frac{\alpha_{k}}{S},k\in[1,...,K]$ (Sensoy et al., 2018).
Given the training set contains $N$ data samples, $D:=\{x_{i},y_{i}\}_{i=1}^{N}$, where $x_{i}$ is the $i^{th}$ data sample and $y_{i}\in[0,1]$ is the corresponding label, 0 is the negative sample and 1 is the positive sample. We further denote $\mathbf{y_{i}}$ as the one-hot encoding label for sample $i$, e.g., $\mathbf{y_{i}}=[1,0]$ for class 0 and $\mathbf{y_{i}}=[0,1]$ for class 1. The original focal loss is designed by Lin et al. (2017), aiming to address the imbalanced class problem and further penalized well-classified samples. The focal loss for binary classification is defined by $FL(p_{t})=-\alpha_{t}(1-p_{t})^{\gamma}\log(p_{t})$ where $p_{t}=p$ if $y_{i}=1$ for $i^{th}$ sample, otherwise $p_{t}=1-p$ with probability output $p$ from the model. Let $\mathbf{P}_{i}$ be a vector that contains the probability of $i^{th}$ sample for both classes from our model output; $p_{i,j}$ is the probability of $i^{th}$ sample belonging to $j^{th}$ class; $K$ is the number of classes, and $\beta_{j}$ is the class weight of $j^{th}$ class. $\gamma$ is the focusing parameter to reduce the loss for well-classified samples, and we fix $\gamma=2$ in this task. We could define Evidential Focal Loss as the following:
@@
Rewriting class probabilities in vector form, the equation (4) can be simplified to (5) by the definition of expectations:
@@
Following the idea of focal loss, we replace the constant term $1$ in the original focal loss function with $\mathbf{y_{i}}$, with the goal to tackle the hard-to-classified samples and reduce the loss of well-classified samples for both classes. Recall that expected probability $\hat{p}_{k}$ for the $k^{th}$ class is $\alpha_{k}/S$, then by the linearity of expectations and the definition of expectations of Dirichlet distribution, we could simply to:
$\mathcal{L}^{cls}_{i}(\theta)=\sum_{j=1}^{K}\boldsymbol{\beta}(y_{ij}-(\alpha_{j}/S))^{2}(\psi(S_{i})-\psi(\alpha_{ij}))$ (6)
where $\psi(\cdot)$ is the digamma function, $y_{ij}$ is the $j^{th}$ class label in the one hot encoding representation $\mathbf{y_{i}}$ and $\boldsymbol{\beta}$ is the class weight vector of length $K$.
To ensure that highly uncertain data samples, referred to as "I do not know" decisions, do not impact the overall data fit and to minimize their associated evidence, we adopt the approach presented in Sensoy et al. (2018). This involves utilizing the Kullback-Leibler (KL) divergence as a regularization term to penalize the unknown predictive distributions, effectively shrinking their influence towards zero. The KL divergence is as same as it is defined in Sensoy et al. (2018).
Finally, our total loss is defined as:
$\mathcal{L}^{total}(\theta)=\sum_{i=1}^{N}\mathcal{L}^{cls}_{i}(\theta)+\lambda_{t}\sum_{i=1}^{N}KL[D(\mathbf{P}_{i}|\alpha_{i})||D(\mathbf{P}_{i}|\mathbf{1})]$ (7)
where $\mathbf{1}$ is an one-vector, $\lambda_{t}$ is the balancing factor between $\mathcal{L}^{cls}$ and the KL divergence loss, and is defined as $\lambda_{t}=min(1.0,t/10)\in[0,1]$, where $t$ is the current number of epochs of training.

### v2

@@ 2.2 Deep Learning for PCa Classification
The use of 3D-CNN models has gained widespread popularity for classifying PCa based on volumetric image data due to their excellent performance. [20] propose a feature fusion 3D-CNN to classify clinically significant PCa using mp-MRI data. They use ADC maps, DWI, and Ktrans 3.0T MR data to enable the model to learn multi-modal information. Inspired by the VGG architecture [79], the model has three VGG-like feature extractors for each image modality, followed by the concatenation between outputs of each extractor and a vector represents the zonal information of the suspicious region. On the test set, the proposed model achieves the area under the receiver operating characteristic (AUC) curve of 0.80 on 140 unseen patients. [19] propose a similar VGG-like 3D-CNN architecture for the same PCa classification task. Different from [20], they only have one model for feature extraction. To obtain the multi-modal information, they stack three images from each of the ADC maps, DWI, Ktrans into one 3-channel image as the input. The model achieves the AUC of 0.84 on the test set.
In [14], a probabilistic approach using mp-MRI data is employed for PCa classification. The authors develop an automated pipeline for the classification of clinically significant PCa using 3.0T DWI images from 427 patients. The pipeline consists of three parts: classification of each DWI slice using the pre-activated ResNet model [80], extraction and selection of first-order statistics from the CNN outputs, and final class label prediction using a random forest classifier. On the test set, the model achieves an AUC of 0.87. While the aforementioned studies may yield favorable AUCs, the reproducibility of the model might be challenging in clinics with limited patient (data) throughput. Recently, [5, 21] address the data-hungry problem by introducing a disentangled representation learning approach (SDNet) to synthesize public 3.0T MRI images into 1.5T MRI images to increase the training data size for centres with limited 1.5T data. Their approach aims to separate the anatomy- and modality-specific features present in images, subsequently merging the 1.5T modality features with the 3.0T anatomical features to generate MRI images resembling those acquired at 1.5T. Finally, a simple 3D-CNN classifier is used for the binary classification of clinically significant PCa. The model outperforms the state-of-the-art performance in PCa classification through domain alignment between different data sources.
While these methods exhibit excellent classification performance, a common limitation is the absence of a confidence score for their predictions, which hinders their interpretability in clinical practice.
@@ 2.3 Uncertainty Estimation
Recent studies in medical imaging have highlighted the detrimental impact of label noise on the performance of modern deep learning models [81]. Conventional regularization techniques such as dropout, batch normalization, weight decay, etc. fail to effectively address this challenge [82, 83]. Methods proposed to mitigate such problem can be broadly categorized into three groups [84]: (1) Robust loss functions and loss adjustments [85, 86, 87] aiming to stabilize the model performance when optimizing its parameters; (2) Sample selection [88, 89, 90] aiming to select a subset of “clean” data from a batch of samples to compute the loss; and (3) Robust architectures [35, 91] aiming to learn the same data by training multiple models with different initialization assess output stability. While these methods inherently handle the noisy label problem, they can not provide explicit uncertainty estimation in terms of confidence in their output. Moreover, the ability of deep learning models to identify irrelevant samples remains limited. For instance, when a model trained on prostate MRIs is presented with a CT scan of the prostate at the time of inference, it is unclear whether the model can provide meaningful predictions or simply indicate a lack of in-domain knowledge and perform a human-in-the-loop analysis instead. In recent years, research has been conducted on uncertainty estimation for deep learning models. [92, 93] develop the dropout neural networks framework to represent the prediction uncertainty of deep learning models, where the dropout layers in the model are formed by Bernoulli distributed random variables. During the test phase, predictive uncertainty is determined by enabling dropout layers and averaging the results over multiple runs, providing a valuable mechanism for uncertainty quantification in model predictions. An alternative method for modeling uncertainty in deep learning models is through the use of evidential neural networks [34], which formulate uncertainty by fitting a Dirichlet distribution - acting as the conjugate prior of the categorical distribution - to the class probabilities acquired from neural networks. This method considers model predictions as multinomial subjective opinions [94] or beliefs [95], which can be further modeled explicitly using subjective logic. The "evidential" approach emphasizes the ability of the model to deliver certain predictions and exhibits superiority compared to the dropout approach [92].
@@
In clinical practice, uncertainty estimation is crucial. By integrating uncertainty information into prediction outcomes, misclassification rates can be significantly reduced. For instance, in radiograph classification task [96], the authors employ the Dempster-Shafer Theory of Evidence [95] and the principles of subjective logic [94] to develop a framework that jointly estimates per-class probabilities and provides predictive uncertainty. This approach has been extended to abdominal ultrasound and brain MR images [97]. In the context of breast cancer classification, [98] apply the evidential neural networks approach [34] to effectively diagnose breast cancer. A similar approach is used for the same task by [99] through the evidence adjustment technique, which focuses on the difference in the risks of uncertain samples from different classes. Consequently, we build upon the work from [34] by adding uncertainty estimation to improve the robustness of the model and the interpretability of predictions.
@@ 3.1 Data
ProstateX Grand Challenge Data (3.0T). The 3.0T data is provided by the International Society of Optics and Photonics in the “ProstateX” challenge [100]. The dataset contains T2-weighted (T2), maximum b-value diffusion, diffusion-weighted imaging (DWI) with apparent diffusion coefficient (ADC) maps, and $K^{trans}$ images of 346 patients undergoing prostate biopsies. T2 images show the anatomical structure of the prostate, and both the ADC maps and Ktrans could further highlight the differences between normal and abnormal (cancer) cells in the MRI scans [101, 102]. We only use 204 of the total 346 patients in this work since these are reserved as training data, and hence they are provided with the spatial location of the suspicious finding, and a binary label indicating whether or not there is cancer. The remaining 142 patients are reserved as the test set and no labels are provided, hence, we exclude those from our work.
@@
Figure 1: Visualization of sample data. 1(a) and 1(b) are the 1.5T T2 and ADC images from KHSC, respectively. Similarly, 1(c) and 1(d) are the 3.0T T2 and ADC images from the “ProstateX” Challenge, respectively.
@@ 3.2 Pre-processing
T2 and ADC sequences from both datasets are $160\times 160\times C$, where $C$ is the total number of slices in the MRI. We resample all 3D data to have the same voxel spacing. To reduce aliasing artifacts, the most common voxel spacing ($0.5\times 0.5\times 3~{}mm^{3}$) is used across all data, and the consine-windowed interpolation is utilized during sampling. We normalize pixel intensities to $[-1,1]$ for all data. For the translation purpose from 3.0T to 1.5T, we further resample all 3D data to $256\times 256\times C$ and split into $C$ 2D gray-scale slices.
@@ 4.1 The Domain Transfer Framework
We adopt the ACL-GAN model [61] for unpaired MR image translation from 3.0T to 1.5T. We have made modifications to the architecture to adapt it to single-channel MRI slices. There are two generators in this model namely $G_{T->S}$ and $G_{S->T}$, $G_{T->S}$ translates the images from the target domain ($X_{T}$) to the source domain ($X_{S}$) given the input $x\in X_{T}$ and a noise vector $z$ sampled from $\mathcal{N}(0,1)$. On the other hand, $G_{S->T}$ performs the reverse process of $G_{T->S}$, translating images from the source domain to the target domain. There are three discriminators, $D_{S},D_{T}$, and $\hat{D}$ in this model. The first two ensure that translated images are in their respective domains by optimizing adversarial losses. The third discriminator, $\hat{D}$, ensures that translated images retain anatomical features in 3.0T by distinguishing the pair (Source, Trans. Source1) and (Source, Trans. Source2), as shown in the bottom of Figure 2. The loss function of ACL-GAN contains four parts [61]. First, the adversarial $\mathcal{L}_{adv}$ loss aims to encourage the translated image, either from source to target domain or from target to source domain, to be in the correct domain. The $\mathcal{L}_{adv}$ is further decomposed to source-domain and target-domain adversarial loss, $\mathcal{L}^{S}_{adv}$ and $\mathcal{L}^{T}_{adv}$, respectively. The $\mathcal{L}^{T}_{adv}$ is given by:
$\mathcal{L}^{T}_{adv}(G_{S->T},D_{T},X_{S},X_{T})$ $=\mathbb{E}_{x_{T}\sim p_{T}}[logD_{T}(x_{T})]$ (1)
$+\mathbb{E}_{\bar{x}_{T}\sim p_{\{\bar{x}_{T}\}}}[log(1-D_{T}(\bar{x}_{T}))]$
where $\bar{x}_{T}=G_{S->T}(x_{S},z_{1})$ and $z_{1}\sim\mathcal{N}(0,1)$. Similarly, the $\mathcal{L}^{S}_{adv}$ is given by:
$\mathcal{L}^{S}_{adv}(G_{T->S},D_{S},\{\bar{x}_{T}\},X_{S})$ $=\mathbb{E}_{x_{S}\sim p_{S}}[logD_{S}(x_{S})]$ (2)
$+(\mathbb{E}_{\hat{x}_{S}\sim p_{\{\hat{x}_{S}\}}}[log(1-D_{S}(\hat{x}_{S}))]$
$+\mathbb{E}_{\tilde{x}_{S}\sim p_{\{\tilde{x}_{S}\}}}[log(1-D_{S}(\tilde{x}_{S}))])/2$
where $\hat{x}_{S}=G_{T->S}(\bar{x}_{T},z_{2})$, $\tilde{x}_{S}=G_{T->S}(x_{S},z_{3}),z_{2},z_{3}\sim\mathcal{N}(0,1)$. Combine these two adversarial losses we get the total adversarial loss for ACL-GAN, $\mathcal{L}_{adv}=\mathcal{L}^{T}_{adv}+\mathcal{L}^{S}_{adv}$.
A problem of $\mathcal{L}_{adv}$ is that this loss can not encourage the translated image to the target domain $\bar{x}_{T}$ is similar to the corresponding source domain image $x_{S}$, as we do not want the model to change the anatomical structure or features as we discussed previously. Hence, the adversarial consistency loss is proposed, which is given by:
$\mathcal{L}_{acl}$ $=\mathbb{E}_{(x_{S},\hat{x}_{S})\sim p_{(X_{S},\{\hat{x}_{S}\})}}[log\hat{D}(x_{S},\hat{x}_{S})]$ (3)
$+\mathbb{E}_{(x_{S},\tilde{x}_{S})\sim p_{(X_{S},\{\tilde{x}_{S}\})}}[log(1-\hat{D}(x_{S},\tilde{x}_{S}))]$
where $\bar{x}_{T}=G_{S->T}(x_{S},z_{1})$, $\hat{x}_{S}=G_{T->S}(\bar{x}_{T},z_{2})$, $\tilde{x}_{S}=G_{T->S}(x_{S},z_{3})$.
Next, $\mathcal{L}_{idt}$ is the identity loss, which encourages generators to perform approximately identity mapping when images in the respective domain are provided, e.g., $x_{S}$ to $G_{T->S}$ or $x_{T}$ to $G_{S->T}$. The $\mathcal{L}_{idt}$ is given by:
$\mathcal{L}_{idt}=\mathbb{E}_{x_{S}\sim p_{S}}[||x_{S}-x_{S}^{idt}||_{1}]+\mathbb{E}_{x_{T}\sim p_{T}}[||x_{T}-x_{T}^{idt}||_{1}]$ (4)
where $E^{z}_{S}:X_{S}\rightarrow Z$ and $E^{z}_{T}:X_{T}\rightarrow Z$ are two noise encoder networks for $G_{S}$ and $G_{T}$, respectively, which map images to noise vectors. $x_{S}^{idt}=G_{T->S}(x_{S},E^{z}_{S}(x_{S}))$ and $x_{T}^{idt}=G_{S->T}(x_{T},E^{z}_{T}(x_{T}))$.
Finally, $\mathcal{L}_{mask}$ is used to force both generators to only modify certain regions of the source image and keep the rest of the areas unchanged. We let generators produce a two-channel image, where the first channel is one of the translated images between the source and target domain (i.e., one-channel gray-scale prostate MRIs), and the second channel is the bounded mask, whose values are between $[0,1]$. The $\mathcal{L}_{mask}$ is given by:
$\mathcal{L}_{mask}$ $=\delta[(\max\{\sum_{k}x_{m}[k]-\delta_{max}\times P_{t},0\})^{2}$ (5)
$+(\max\{\delta_{min}\times P_{t}-\sum_{k}x_{m}[k],0\})^{2}]$
$+\sum_{k}\frac{1}{|x_{m}[k]-0.5|+\epsilon}$
where $\delta$, $\delta_{max}$ and $\delta_{min}$ are hyper-parameters for controlling the size of masks, $x_{m}[k]$ is the k-th pixel of the mask and $P_{t}$ is the total number of pixels in an image. The $\epsilon$ is a very small value to avoid dividing by zero. The first term of this loss controls the size of the mask. It encourages the generator to perform sufficient modifications while preserving the background information. Here, $\delta_{max}$ and $\delta_{min}$ are the maximum and minimum proportions of the foreground in the mask, i.e., the region we want to modify. The last term of this loss encourages the mask to be binary, either 0 or 1 to segment the foreground and background of the input image.
Aggregating all loss terms together, we have:
@@
where $\lambda_{acl},\lambda_{idt},\lambda_{mask}$ are the weighting factors for $\mathcal{L}_{acl},\mathcal{L}_{idt},\mathcal{L}_{mask}$, respectively.
@@ 4.2.1 Classifier architectures
The traditional CNN approach is used for the clinically significant PCa binary classification task. Specifically, we explore three different model architectures for combinations of T2 and ADC patches as the classifier (Top row of Figure 2). The first architecture, called the multi-stream CNN (M.S. MpMRI), treats T2 and ADC patches as separate inputs, as shown in Figure 3. The model takes 3D patches of T2 and ADC as parallel inputs, which are then processed by the same feature extractor (i.e., weights are shared) to extract deep semantic representations. The output representations of T2 and ADC are then concatenated channel-wise and fed into another convolutional layer followed by a fully connected layer to produce the class probabilities.
In the second architecture, we adopt two different ways to combine ADC and T2 patches into a single input for the network. First, we stack cropped 3D patches of T2 and ADC along the channel axis, resulting in an input data size of $64\times 64\times 6$. Alternatively, we consider only the located suspicious slice $i_{s}$ for both T2 and ADC, and stack them along the channel axis to obtain the input data size of $64\times 64\times 2$. The model architecture for both combinations is similar to Figure 3, where there is only one branch and no concatenation afterward. We refer to the model with input size of $64\times 64\times 6$ (resp. input size $64\times 64\times 2$) as Vol. MpMRI (resp. MpMRI).
Lastly, we use only 3D T2 patches as input to match with the previous work [5]. The model architecture is as same as the one for MpMRI, and we refer to this model as “T2-only”.
@@
We also explore with the co-teaching framework [35] to combat the potential noisy label issue to improve the classification performance. This framework can be directly used as our classifier in a plug-and-play manner. The key distinction between the co-teaching framework and the conventional training paradigm lies in the fact that co-teaching only influences the number of samples in the batch that contribute to the loss calculation. We follow the exact setup in the original paper, and detailed hyperparameter settings can be found in Section 5. The backbone model we selected for this framework is the “MpMRI” architecture, which we depicted previously.
@@ 4.2.2 Evidential Focal Loss
Dataset filtering during the training phase could reduce the effect of the noisy label on the deep model. Followed by [96], the process of uncertainty-based filtering is shown at the top of Figure 2: Firstly, we calculate the uncertainty value for each sample in the training set. We then remove a portion of the training samples that exhibit high predictive uncertainty. Finally, we retrain the model using the remaining “clean” training data.
Following the previous work [34], we extend and combine the idea of subjective logic [94] with the focal loss [36] for the clinically significant PCa binary classification task. In the context of the Theory of Evidence, a belief mass is assigned to individual attributes, e.g., the possible class label of a specific data sample. The belief mass is generally calculated from the evidence collected from the observed data [95]. Let $K$ be the number of classes, and $b_{k}\geq 0,k\in[1,K]$ is the belief mass for class $k$ and $u\geq 0$ is the overall uncertainty measure. Let $e_{k}\geq 0$ be the evidence computed for $k^{th}$ class, then the belief $b_{k}$ and the uncertainty $u$ are computed as follows:
@@
where $S=\sum_{i=1}^{K}(e_{i}+1)$. For our binary task ($K=2$), we can further simplify Equation (7) to $b_{0}=\frac{e_{0}}{e_{0}+e_{1}+2},b_{1}=\frac{e_{1}}{e_{0}+e_{1}+2}$, and $u=\frac{2}{e_{0}+e_{1}+2}$. The belief mass assignment, e.g., subjective opinion, corresponds to the Dirichlet distribution with parameters $\alpha_{k}=e_{k}+1$, and $S=\sum_{i=1}^{K}\alpha_{k}$ is the Dirichlet strength. The expected probability for the $k^{th}$ class is calculated by the mean with the associate Dirichlet distribution, e.g., $\hat{p}_{k}=\frac{\alpha_{k}}{S},k\in[1,...,K]$ [34].
Now, we formally define our proposed evidential focal loss. Given the training set contains $N$ data samples, $D:=\{x_{i},y_{i}\}_{i=1}^{N}$, where $x_{i}$ is the $i^{th}$ data sample and $y_{i}\in[0,1]$ is the corresponding label, 0 is the negative sample and 1 is the positive sample. We further denote $\mathbf{y_{i}}$ as the one-hot encoding label for sample $i$, e.g., $\mathbf{y_{i}}=[1,0]$ for class 0 and $\mathbf{y_{i}}=[0,1]$ for class 1. The focal loss [36] for binary classification is defined by $FL(p_{t})=-w_{t}(1-p_{t})^{\gamma}\log(p_{t})$ where $p_{t}=p$ if $y_{i}=1$ for $i^{th}$ sample, otherwise $p_{t}=1-p$ with probability output $p$ from the model, and $w$ is the class weight. Let $\mathbf{P}_{i}$ be a vector that contains the probability of $i^{th}$ sample for both classes from our model output; $p_{ij}$ is the probability of $i^{th}$ sample belonging to $j^{th}$ class; $K$ is the number of classes, and $\beta_{j}$ is the class weight of $j^{th}$ class. $\gamma$ is the focusing parameter to reduce the loss for well-classified samples, and we fix $\gamma=2$ in this task. We could define Evidential Focal Loss as the following:
@@
Here, we denote $\mathcal{L}^{cls}_{i}(\theta)$ as the classification loss for a single sample $i$, $\alpha_{i}$ is the belief mass $\alpha$ for $i^{th}$ sample for all classes and $\alpha_{ij}$ is the belief mass $\alpha$ for $i^{th}$ sample and $j^{th}$ class. $\mathcal{B}(\cdot)$ is the multinomial beta function and $\frac{1}{\mathcal{B}(\alpha_{i})}\prod_{j=1}^{K}p_{ij}^{\alpha_{ij}-1}$ is the probability density function of Dirichlet distribution. Rewriting class probabilities in vector form, the equation (8) can be simplified to (9) by the definition of expectations:
@@
Different from the original focal loss, we replace the constant term $1$ with $\mathbf{y_{i}}$, to tackle the hard-to-classified samples and reduce the loss of well-classified samples in both classes. Recall that expected probability $\hat{p}_{k}$ for the $k^{th}$ class is $\alpha_{k}/S$, then by the linearity of expectations and the definition of expectations of Dirichlet distribution, we have:
$\mathcal{L}^{cls}_{i}(\theta)=\sum_{j=1}^{K}{\beta_{j}}(y_{ij}-(\alpha_{j}/S))^{2}(\psi(S_{i})-\psi(\alpha_{ij}))$ (10)
where $\psi(\cdot)$ is the digamma function, $y_{ij}$ is the $j^{th}$ class label in the one hot encoding representation $\mathbf{y_{i}}$ and $\beta_{j}$ is the class weight vector for class $j$.
To ensure that highly uncertain data samples, referred to as "I do not know" decisions, do not impact the overall data fit and to minimize their associated evidence, we adopt the Kullback-Leibler (KL) loss as the regularization term to penalize the unknown predictive distributions, as done in [34]. Combining $\mathcal{L}^{cls}$ and the KL loss yields our final evidential focal loss for uncertainty-aware classification:
$\mathcal{L}^{efl}(\theta)=\sum_{i=1}^{N}\mathcal{L}^{cls}_{i}(\theta)+\lambda_{t}\sum_{i=1}^{N}KL[D(\mathbf{P}_{i}|\alpha_{i})||D(\mathbf{P}_{i}|\mathbf{1})]$ (11)
where $\mathbf{1}$ is an one-vector, $D(\mathbf{P}_{i}|\mathbf{1})$ is the uniform Dirichlet distribution. $\lambda_{t}$ is the weighting factor of the KL divergence loss and is defined as $\lambda_{t}=min(1.0,t/10)\in[0,1]$, where $t$ is the current number of epochs of training.